OpenStudy (anonymous):

What is a counterexample for the conjecture? Conjecture: Any number that is divisible by 4 is also divisible by 8. A.) 24 B.) 40 C.) 12 D.)26

OpenStudy (anonymous):

For example 4 :) Can you divided 4 on 8 and have a number in N :)

OpenStudy (anonymous):

Im sorry i do not understand

OpenStudy (anonymous):

12 is a counter

OpenStudy (anonymous):

any number where N/4 = an ODD number will give you N/8= a decimal of X.5 . Examples: 12 (12/8= 1.5), 20 (20/8= 2.5), and 44 (44/8=5.5)